# The Wavefront Set: Bounds for the Langlands parameter

### v1

@@ 4. Classical groups (after Lust-Stevens [LS])
where $s$ runs over the conjugacy classes of semisimple elements in ${{\mathcal{G}}^{*}}^{F}$. Since $Z({\mathcal{G}})$ is connected, the rational semisimple conjugacy classes in ${\mathcal{G}}^{*}$ are in one-to-one correspondence with the geometric $F$-stable semisimple conjugacy classes in ${\mathcal{G}}^{*}$ [Car, Proposition 3.7.3].
The above decomposition is realized as follows. By [Car, Theorem 4.4.6], there exists a natural one-to-one correspondence between geometric conjugacy classes of pairs $(T,\theta)$, where ${\mathcal{T}}$ is a maximal $F$-stable torus in ${\mathcal{G}}$ and $\theta$ is a character of $T$, and $F$-stable semisimple conjugacy classes in ${\mathcal{G}}^{*}$. Let $s\in{\mathcal{T}}^{*}\subset{\mathcal{G}}^{*}$, then $\mathsf{Irr}_{s}{\mathcal{G}}^{F}$ consist of all the constituents of the induced representations $R_{{\mathcal{T}}}^{{\mathcal{G}}}(\theta)$ where…
@@
$\Psi_{s}^{{\mathcal{G}}}:\mathsf{Irr}_{s}{\mathcal{G}}^{F}\longleftrightarrow\mathsf{Irr}_{s}({\mathcal{G}}_{s}^{F},1),$
where ${\mathcal{G}}_{s}$ is the endoscopic group of ${\mathcal{G}}$ with dual ${\mathcal{G}}^{*}_{s}:=Z_{{\mathcal{G}}^{*}}(s)$. To have cuspidal representations, we need $Z({\mathcal{G}}^{*})^{0}$ and $Z({\mathcal{G}}_{s}^{*})^{0}$ to have the same $F$-rank. In this case, $\Psi_{s}^{{\mathcal{G}}}$ gives a correspondence of cuspidal represenattions. In other words, the set of cuspidal ${\mathcal{G}}^{F}$-representations are in one-to-one correspondence with ${\mathcal{G}}^{*}$-pairs $(s,\tau)$ where $s\in{{\mathcal{G}}^{*}}^{F}$ is semisimple of equal rank and $\tau$ is an cuspidal unipotent representation of ${\mathcal{G}}_{s}^{F}$.
@@
where the product runs over all irreducible self-dual monic polynomials over ${\mathbb{F}}_{q}$ and nonnegative integers $a_{P}$ such that $\sum_{P}a_{P}\deg(P)=\dim V$, where $V$ is the defining representation of ${\mathcal{G}}^{*}$. In order for the semisimple class to parametrize a cuspidal representation, one needs ([LS]):
@@
    if we set $a_{\pm}=a_{x\mp 1}$, there are some conditions on $a_{\pm}$ depending on the type of the group, see [LS, §7.2].
@@
4.1. Nilpotent Langlands parameter
@@
where $K$ is the normalizer of a maximal parahoric subgroup with reductive quotient $\bar{K}={{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}$ over ${\mathbb{F}}_{q}$, and $\tau_{1}\boxtimes\tau_{2}$ is a cuspidal representation of ${{\mathcal{G}}}^{\circ}_{N_{1},N_{2}}({\mathbb{F}}_{q})$. The explicit cases are as follows [LS, §2], for more details about the notation, see loc. cit.:
@@
Let $(m_{\pm}^{(i)},m_{P}^{(i)},k_{P}^{(i)})$, $i=1,2$ be the collections of integers that characterizes $\tau_{i}$. The nilpotent Langlands parameter of $\pi$ is a nilpotent orbit ${\mathbb{O}}^{\vee}_{\pi}$ in $\mathfrak{g}^{\vee}$, which is either $gl(n,{\mathbb{C}})$, $sp(2n,{\mathbb{C}})$ or $so(m,{\mathbb{C}})$. Via the Jordan normal form, ${\mathbb{O}}^{\vee}_{\pi}$ is parametrized by a partition $\lambda$ of $n$, $2n$, or $m$, respectively. If $\mathfrak{g}^{\vee}=so(m,{\mathbb{C}})$, then the partition is orthogonal, that is each even part appears with even multiplicity. If $\mathfrak{g}^{\vee}=sp(2n,{\mathbb{C}})$, then the partition is symplectic, that is each odd part appears with even multiplicity.
Lust and Stevens [LS, §8] attach to $\pi$ the nilpotent orbit ${\mathbb{O}}^{\vee}_{\pi}$ given by the partition obtained by the following procedure. For every integer $\alpha$ denote by $[\alpha]$ the string $\{\alpha,\alpha-2,\alpha-4,\dots,1\}$ if $\alpha>0$ is odd, $\{\alpha,\alpha-2,\alpha-4,\dots,2\}$ if $\alpha>0$ is even, and $[\alpha]=\emptyset$ if $\alpha\leq 0$. Start with an empty partition and
@@ 4.2. Symplectic groups
If $\alpha,\beta$ are two partitions of $m,n$, respectively, the sum $\alpha+\beta$ is a partition of $m+n$ given as follows. Suppose $\alpha=(\alpha_{1},\alpha_{2},\dots,\alpha_{\ell})$ and $\beta=(\beta_{1},\beta_{2},\dots,\beta_{\ell})$ with $0\leq\alpha_{1}\leq\alpha_{2}\leq\dots\leq\alpha_{\ell}$, $0\leq\beta_{1}\leq\beta_{2}\leq\dots\leq\beta_{e}ll$. (By allowing entries of 00, we may assume there is the same number of parts in $\alpha$ and $\beta$.) Then $\alpha+\beta=(\alpha_{1}+\beta_{1},\alpha_{2}+\beta_{2},\dots,\alpha_{\ell}+\beta_{\ell}).$
This is operation on partitions plays a key role in the explicit computation of $j$-induction for representations of the Weyl group of classical types.
@@
$\sum_{\deg P>1}k_{p}a_{p}^{(i)}=n_{i},\quad a_{P}^{(i)}=\frac{{m_{P}^{(i)}}^{2}+m_{P}^{(i)}}{2}.$
To compute the wavefront set, we need to compute the $j$-induction on each factor. To simplify the notation, let us drop temporarily the indices $i$. The calculation amounts to determining the irreducible $W_{n}$-representation given
@@
Using the formulae for $j$-induction [Lu3], This equals
@@
Now suppose both $n_{1},n_{2}>0$. The key observation is that for all $k$, $m^{(1)}\geq m^{(2)}$,
@@
Using the formulae for $j$-induction from [Lu3], we see this $j$-induced representation is the bipartition
@@
Let $\lambda_{\pm}^{\vee}$ be the orthogonal partition for the dual nilpotent orbit given by the Lust-Stevens algorithm for the polynomials $P$ with degree equal to $1$. That is, $\lambda_{\pm}^{v}ee$ has an expression like in Proposition 4.5(1). Let $\mu_{\pm}$ be the symplectic partition for the wavefront set again for the $\deg P>1$ part as in Proposition 4.5(2).
@@ 4.4. Special orthogonal groups
Let $\pi$ be a depth-zero supercuspidal $SO(m,\mathsf{k})$-representation. Suppose there are no endoscopic unitary factors, i.e., no polynomials $P$ with $\deg P>1$ appear. Then
@@ 4.5. Ramified unitary groups
For the ramified unitary groups, the finite group ${{\mathcal{G}}}^{\circ}_{N_{1},N_{2}}({\mathbb{F}}_{q})$ on which the cuspidal representation lives is a product of a symplectic and special orthogonal group. For notational purposes, assume ${{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}({\mathbb{F}}_{q})=Sp(2n_{1},{\mathbb{F}}_{q})\times SO(n_{2}+n_{2}^{an},n_{2},{\mathbb{F}}_{q})$. We only record the analogue of Propositions 4.5 and 4.7, the proof is simply a combination of the symplectic and orthogonal cases discussed before.
@@
Let $\pi$ be a depth-zero supercuspidal representation of a ramified unitary group induced from ${{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}$ as above. Suppose there are no endoscopic unitary factors, i.e., no polynomials $P$ with $\deg P>1$ appear. Then

### v2

@@ 4. Depth-zero representations of classical groups
where $s$ runs over the conjugacy classes of semisimple elements in ${{\mathcal{G}}^{*}}^{F}$. The piece $\mathsf{Irr}_{1}{\mathcal{G}}^{F}$ is the set of irreducible unipotent representations. Since $Z({\mathcal{G}})$ is connected, the rational semisimple conjugacy classes in ${\mathcal{G}}^{*}$ are in one-to-one correspondence with the geometric $F$-stable semisimple conjugacy classes in ${\mathcal{G}}^{*}$ [Car, Proposition 3.7.3].
The above decomposition is realized as follows. By [Car, Theorem 4.4.6], there exists a natural one-to-one correspondence between geometric conjugacy classes of pairs $(T,\theta)$, where ${\mathcal{T}}$ is a maximal $F$-stable torus in ${\mathcal{G}}$ and $\theta$ is a character of $T$, and $F$-stable semisimple conjugacy classes in ${\mathcal{G}}^{*}$. Let $s\in{\mathcal{T}}^{*}\subset{\mathcal{G}}^{*}$, then $\mathsf{Irr}_{s}{\mathcal{G}}^{F}$ consists of all the constituents of the induced representations $R_{{\mathcal{T}}}^{{\mathcal{G}}}(s)$ where we think of $s$ as a character of $T$.
@@
$\Psi_{s}^{{\mathcal{G}}}:\mathsf{Irr}_{s}{\mathcal{G}}^{F}\longleftrightarrow\mathsf{Irr}_{1}({\mathcal{G}}_{s}^{F}),$
where ${\mathcal{G}}_{s}$ is the endoscopic group of ${\mathcal{G}}$ with dual ${\mathcal{G}}^{*}_{s}:=Z_{{\mathcal{G}}^{*}}(s)$. To have cuspidal representations, we need $Z({\mathcal{G}}^{*})^{0}$ and $Z({\mathcal{G}}_{s}^{*})^{0}$ to have the same $F$-rank. In this case, $\Psi_{s}^{{\mathcal{G}}}$ gives a correspondence of cuspidal represenattions. In other words, the set of cuspidal ${\mathcal{G}}^{F}$-representations are in one-to-one correspondence with ${\mathcal{G}}^{*}$-pairs $(s,\tau)$ where $s\in{{\mathcal{G}}^{*}}^{F}$ is semisimple of equal rank and $\tau$ is a cuspidal unipotent representation of ${\mathcal{G}}_{s}^{F}$.
@@
where the product runs over all irreducible self-dual monic polynomials over ${\mathbb{F}}_{q}$ and nonnegative integers $a_{P}$ such that $\sum_{P}a_{P}\deg(P)=\dim V$, where $V$ is the defining $\mathbb{F}_{q}$-representation of ${\mathcal{G}}^{*}$. In order for the semisimple class to parametrize a cuspidal representation, one needs:
@@
    if we set $a_{\pm}=a_{x\mp 1}$, there are some conditions on $a_{\pm}$ depending on the type of the group, see [LuSt, §7.2].
@@
4.1. Nilpotent Langlands parameters
@@
where $K$ is the normalizer of a maximal parahoric subgroup with reductive quotient $\bar{K}={{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}$ over ${\mathbb{F}}_{q}$, and $\tau_{1}\boxtimes\tau_{2}$ is a cuspidal representation of ${{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}({\mathbb{F}}_{q})$. The explicit cases are as follows [LuSt, §2], for more details about the notation, see loc. cit.:
@@
Let $(m_{\pm}^{(i)},m_{P}^{(i)},k_{P}^{(i)})$, $i=1,2$ be the collections of integers that characterizes $\tau_{i}$. The nilpotent Langlands parameter of $\pi$ is a nilpotent orbit ${\mathbb{O}}^{\vee}_{\pi}$ in $\mathfrak{g}^{\vee}$, which is either $gl(n,{\mathbb{C}})$, $sp(2n,{\mathbb{C}})$ or $so(m,{\mathbb{C}})$. Via the Jordan normal form, ${\mathbb{O}}^{\vee}_{\pi}$ is parametrized by a partition $\lambda$ of $n$, $2n$, or $m$, respectively. If $\mathfrak{g}^{\vee}=so(m,{\mathbb{C}})$, then the partition is orthogonal: each even part appears with even multiplicity. If $\mathfrak{g}^{\vee}=sp(2n,{\mathbb{C}})$, then the partition is symplectic: each odd part appears with even multiplicity.
Lust and Stevens [LuSt, §8] attach to $\pi$ the nilpotent orbit ${\mathbb{O}}^{\vee}_{\pi}$ given by the partition obtained by the following procedure. For every integer $\alpha$ denote by $[\alpha]$ the string $\{\alpha,\alpha-2,\alpha-4,\dots,1\}$ if $\alpha>0$ is odd, $\{\alpha,\alpha-2,\alpha-4,\dots,2\}$ if $\alpha>0$ is even, and $[\alpha]=\emptyset$ if $\alpha\leq 0$. Start with an empty partition and
@@ 4.2. Symplectic groups
If $\alpha,\beta$ are two partitions of $m,n$, respectively, the sum $\alpha+\beta$ is a partition of $m+n$ given as follows. Suppose $\alpha=(\alpha_{1},\alpha_{2},\dots,\alpha_{\ell})$ and $\beta=(\beta_{1},\beta_{2},\dots,\beta_{\ell})$ with $0\leq\alpha_{1}\leq\alpha_{2}\leq\dots\leq\alpha_{\ell}$, $0\leq\beta_{1}\leq\beta_{2}\leq\dots\leq\beta_{\ell}$. (By allowing entries of 00, we may assume there is the same number of parts in $\alpha$ and $\beta$.) Then $\alpha+\beta=(\alpha_{1}+\beta_{1},\alpha_{2}+\beta_{2},\dots,\alpha_{\ell}+\beta_{\ell}).$
This operation on partitions plays a key role in the explicit computation of $j$-induction for representations of the Weyl group of classical types.
@@
$\sum_{\deg P>1}k_{p}a_{p}^{(i)}=n_{i},\quad a_{P}^{(i)}=\frac{{m_{P}^{(i)2}}+m_{P}^{(i)}}{2}.$
To compute the wavefront set, we need to compute the $j$-induction on each factor. To simplify the notation, let us drop temporarily the indices $i$. Let $W_{n}$ be the Weyl group of $G^{\vee}$. Then, the calculation amounts to determining the irreducible $W_{n}$-representation given by
@@
Using the formulae for $j$-induction [Lu3, §3.1,§4.4], this equals
@@
Now suppose both $n_{1},n_{2}>0$. The key observation is that for all $k$ and $m^{(1)}\geq m^{(2)}$,
@@
Using the formulae for $j$-induction from [Lu3, §5], we see this $j$-induced representation is the bipartition
@@
Let $\lambda_{\pm}^{\vee}$ be the orthogonal partition for the dual nilpotent orbit given by the Lust-Stevens algorithm for the polynomials $P$ with degree equal to $1$. That is, $\lambda_{\pm}^{\vee}$ has an expression like in Proposition 4.5(1). Let $\mu_{\pm}$ be the symplectic partition for the wavefront set again for the $\deg P>1$ part as in Proposition 4.5(2).
@@ 4.4. Special orthogonal groups
Let $\pi$ be a depth-zero supercuspidal $SO(m,\mathsf{k})$-representation. Suppose there are no endoscopic unitary factors, i.e., no polynomials $P$ with $\deg P>1$ appear. Then,
@@ 4.5. Ramified unitary groups
For the ramified unitary groups, the finite group ${{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}({\mathbb{F}}_{q})$ on which the cuspidal representation lives is a product of a symplectic and special orthogonal group. For notational purposes, assume ${{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}({\mathbb{F}}_{q})=Sp(2n_{1},{\mathbb{F}}_{q})\times SO(n_{2}+n_{2}^{an},n_{2},{\mathbb{F}}_{q})$. We only record the analogue of Propositions 4.5 and 4.7. The proof is simply a combination of the symplectic and orthogonal cases discussed before.
@@
Let $\pi$ be a depth-zero supercuspidal representation of a ramified unitary group induced from ${{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}$ as above. Suppose there are no endoscopic unitary factors, i.e., no polynomials $P$ with $\deg P>1$ appear. Then,